OpenStudy (anonymous):

Use Elimination to solve -3x+2y=10 -2x-y=-5 and what is the process to solve it?

OpenStudy (saifoo.khan):

multiply 2nd eq by two.

OpenStudy (anonymous):

you can multiply the second eqation by 2 which makes the y's cancel. this makes it possible to solve for x and i think u kno where to go from there :)

OpenStudy (anonymous):

(-2x-y=-5)2 then you will have -4x-2y=-10 add eq 1st to 2nd you will get -7x = 0 means x equal to 0 then you subs it in either eq you will get y = 5

OpenStudy (toxicsugar22):

x=-1, y=7/2

OpenStudy (toxicsugar22):

x=0,y=5